OpenStudy (anonymous):

find the value(s) of c guaranteed by the mean value theorem for integrals for the function over the given interval. f(x)=9/x^3 [1,3]

OpenStudy (anonymous):

ok that was wrong, let me start again \[f(3)=\frac{9}{3^3}=\frac{1}{3}\] \[f(1)=9\] \[\frac{f(3)-f(1)}{3-1}=\frac{\frac{1}{3}-9}{2}= -\frac{13}{3}\]

OpenStudy (anonymous):

then find \[f'(x)=-\frac{27}{x^4}\] and finally solve \[-\frac{27}{x^4}=-\frac{13}{3}\] for x

OpenStudy (anonymous):

this is what I got F(b)-F(a)=4 9/c^3=2 ....am I right so far?.....how did u get 1/3?

OpenStudy (anonymous):

\[f(3)=\frac{9}{3^3}=\frac{1}{3}\] did i make a mistake it is late

OpenStudy (anonymous):

oh okay, hold on let me check

OpenStudy (anonymous):

you have \[f(x)=\frac{9}{x^3}\] so i get \[f(3)=\frac{1}{3}\] and \[f(1)=9\]

OpenStudy (anonymous):

also \[f'(x)=-\frac{27}{x^4}\]

OpenStudy (anonymous):

how did u get that for f'?

OpenStudy (anonymous):

in general can you tell me the steps to finding c?

OpenStudy (anonymous):

\[f(x)=\frac{9}{x^3}=9x^{-3}\] \[f'(x)=-27x^{-4}=-\frac{27}{x^4}\]

OpenStudy (anonymous):

ok lets go slow your job is to find \[\frac{f(b)-f(a)}{b-a}\] which of course is a number i got \[-\frac{13}{3}\] as the number

OpenStudy (anonymous):

then i found \[f'(x)=-\frac{27}{x^4}\]

OpenStudy (anonymous):

which means \[f'(c)=-\frac{27}{c^4}\]

OpenStudy (anonymous):

last job is to set them equal and solve for "c"

OpenStudy (anonymous):

\[-\frac{27}{c^4}=-\frac{13}{3}\] solve for c

OpenStudy (anonymous):

\[13c^4=81\] \[c^4=\frac{81}{13}\] \[c=\sqrt[4]{\frac{81}{13}}\]

OpenStudy (anonymous):

or \[c=\frac{3}{\sqrt[4]{13}}\]

OpenStudy (anonymous):

thanks for the clarification